# Learning to rank quantum circuits for hardware-optimized performance enhancement

### v1

@@ I Introduction
Extrapolating from hardware-level proxies to system-level performance is complicated by the fact that most common scalable QCVV approaches are by design insensitive to the myriad sub-structures, or contexts, that can appear within quantum circuits and which can have important impacts on the performance. This is in part due to the historic use of these metrics in analyses of fault tolerance, in which errors are typically (and speciously) assumed to be statistically independent. For example, randomized benchmarking (RB) is relatively efficient to implement and easy to interpret, yet it masks coherent errors exhibiting correlations in time over many gates in a circuit [10, 11]. Similarly, crosstalk may change both the magnitude and the nature of the error of a specific gate depending on the existence or absence of simultaneous gates on neighboring qubits [11, 12] - contexts that are ignored over averaged over in standard gate-level QCVV protocols. In particular, errors that occur during idling periods are in general poorly captured by most standard methods, both at the single qubit level where $T_{2}$ effects are relevant, and at the multi-qubit level where quantum crosstalk can lead to an unwanted unitary evolution between neighboring qubits [13, 14, 15]. Furthermore, the overall effect of an error on the measured fidelity of a quantum circuit will be dependent on the structure of the circuit; random circuits will generally be less sensitive to over-rotation gate errors due to self-cancellations of the random excess angles [16], whereas circuits with a repetitive structure, such as Trotter circuits, will be very sensitive to these types of errors [8, 10, 17]. The aggregate effect of these various circumstances - all of which are routinely encountered in the execution of real algorithms on hardware - tends to cause QCVV protocols to only be loosely predictive of algorithmic performance [9, 14].
Here, we focus on the concrete problem of layout selection within a single processor: the mapping of virtual or abstract qubits in a compiled circuit to physical device qubits in a manner that maximizes circuit fidelity. Previous work has examined the topic of layout selection via various methods [18, 19, 20, 21, 22, 23], but has not presented a comprehensive ranking procedure based on measurements of real circuits executed on hardware. This omission ignores the essential contextual information that is known to cause divergence between proxy measure predictions and achievable performance.
@@
In this work, we develop a new form of predictive QCVV for layout selection for arbitrary target circuits based on a heuristic, machine learning solution. We report direct measurements on an IBM quantum computer revealing that different layouts of individual target circuits exhibit up to a factor of $\times1$ variation in the ratio of the maximum achieved to median Hellinger fidelity, highlighting the importance of this task. We develop a procedure to rank the expected relative performance of logically equivalent layouts for arbitrary target circuits within a quantum processor, drawing from a large body of work in machine learning where this general problem is known as “Learning to Rank” [24]. In so doing, this method prioritizes the relative performance between logically equivalent permutations of a user-defined target circuit over accurate estimates of absolute performance, and omits the reliance on human-interpretable proxy measures. Our method trains a circuit-score model over a dataset of circuits exhibiting diverse structural characteristics, and parameterizes the target-circuit score in a physics-based phenomenological manner. The model then assigns to each layout of the target circuit a numerical score used to quantify relative expected performance (a higher score corresponds to a higher expected performance). We test the efficacy of the ranking procedure using direct fidelity measurements of the different layouts on hardware, and compare our learning-to-rank method against alternative layout-selection approaches. We also compare five different candidate loss functions and quantitatively demonstrate that the listwise Rank MSE loss function performs best across multiple metrics. For the best model, measurements demonstrate a reduction in selection error – the relative loss in Hellinger fidelity due to sub-optimal layout selection – by 3.2×3.2$\times$3.2 × (1.8×1.8$\times$1.8 ×) when compared to a random and a baseline error-informed layout selection strategy, respectively, validating the utility of this approach.
The rest of the article is organized as follows. In Sec. II we briefly review the problem of layout selection. In Sec. III we introduce our approach, and then perform experiments on real quantum hardware in Sec. IV, before concluding with a discussion in Sec. V. Further technical details are included in Appendix A and B.
II The layout-selection problem
@@
III Circuit Ranking for Layout Selection
@@
We now present an overview of a practical and efficient machine-learning approach for ranking-based layout selection. We use a score-based approach for ranking circuit performance. The circuit score $S$ is a function that maps scheduled circuits to a real-valued numerical score: higher scores correspond to higher estimated fidelities. In a good ranking procedure, ordering layouts by circuit score should lead to the same ordinal ranking among circuits as a chosen performance metric derived from measurements on hardware. In this work we take this measured quantity to be the Hellinger fidelity $H_{F}$ between the ideal bitstring distribution and the empirically observed bitstring distribution.
@@
$H_{F}(C_{1})\lesssim H_{F}(C_{2})\quad\text{if and only if}\quad S(C_{1})\lesssim S(C_{2})\,.$ (1)
Our objective is now to learn score functions which approximately obey Eq. 1.
@@
The loss function provides a continuous and differentiable way to penalize deviations from Eq. 1. We consider multiple loss functions, summarized in Table 1, reflecting the fact that there does not appear to be a clear preferred choice in the Learning-to-Rank literature. These can be grouped into three categories: pointwise, pairwise, and listwise. Pointwise approaches do not involve any comparisons, pairwise approaches involve comparisons between pairs of layouts, and listwise approaches involve comparisons across a set of all possible layouts for a given initial circuit. Details for the various loss functions considered are presented in Appendix B.
@@
Score MSE
pointwise
Minimize difference between score and Hellinger fidelity
Pearson
listwise
Maximize correlation between score and Hellinger fidelity
Soft Spearman
listwise
Maximize correlation between score and Hellinger fidelity
Rank MSE
listwise
Minimize difference of true and predicted rank
NLL
listwise
Maximize model probability of observed top-$K$ layouts
@@
A summary of the machine learning pipeline for layout selection is presented in Fig. 3. In order to train the ML model a dataset of circuits is first generated. This is illustrated in Fig. 3(a), where each layer represents a batch of layouts transpiled to the hardware for a different circuit. Each layout is then executed on the quantum processing unit (QPU) and the Hellinger fidelity is computed for all layouts by comparing the ideal bitstring distribution against the empirical distribution. Together, these batches of executed layouts and their fidelities form a dataset of labeled data. In the training phase, illustrated in Fig. 3(b), the batches of layouts in the dataset are ranked and the model iteratively updated. For each batch, the current state of the model produces a score ranking that is compared against the true ranking derived from the Hellinger fidelity. Mismatches in the two rankings are penalized by the differentiable loss $\mathcal{L}(\theta)$. The model parameters $\theta$ are updated through multiple iterations of a gradient-based algorithm. In the inference phase, illustrated in Fig. 3(c), the goal is to use the trained model to predict a single layout for execution on the QPU. The input is a single transpiled circuit and the most up-to-date backend information (such as gate errors and $T_{1}$ times). A set of all valid layouts is generated, and a score $S$ is assigned to each layout by the already trained model. The layouts are ranked according to the circuit score, and the circuit with the largest score is chosen for execution on the device.
IV Experimental Results
@@
We benchmark our approach against both random selection and selection via the Mapomatic package [30]. Random selection is the default approach when utilizing a standard compiler. Many informed users employ Mapomatic, a layout selection method that estimates circuit fidelity by aggregating the backend-estimated fidelities of each individual operation (i.e., gate errors, measurement errors, etc).  [^2] Note that Mapomatic is the name of both the layout selection algorithm as well as the Python package used to implement that algorithm - here, we will use Mapomatic to specifically refer to the layout selection algorithm. Mapomatic also uses a scoring function to rank layouts; the key points of difference are in the choice of score function and the fact that in the models developed here, the parameters are learned by training the model over a dataset containing a diverse range of circuit contexts.
@@
Table 2: Model performance. The performance of the circuit score model is measured by multiple evaluative metrics. Each row corresponds to a circuit score model trained on a different loss function, averaged over 10 random seeds. Uncertainties correspond to 1 standard error. For comparison, the results for random layout selection and Mapomatic are also shown. The win rate of each method against Mapomatic (MM) is shown in the final column.
Model Loss Median normed rank Median rank Selection error (%) Top-1 Acc (%) Win Rate vs MM (%)
@@
Rank MSE $(d=1)$ 0.13275(39) 3.70(15) 10.603(97) 22.41(41) 67.19(18)
Rank MSE $(d=2)$ 0.13275(39) 3.70(15) 10.603(97) 22.41(41) 67.19(18)
Rank MSE $(d=3)$ 0.13304(29) 3.80(13) 10.638(78) 22.52(46) 67.03(18)
@@ Appendix A Phenomenological circuit score
[…]
@@
$F(\rho(0),\rho(t))=|\alpha|^{2}$ $+\left(|\alpha|^{4}+|\beta|^{4}-|\alpha|^{2}\right)e^{-\Gamma_{1}t}$ (10)
$+2|\alpha|^{2}|\beta|^{2}e^{-\Gamma_{2}t}\,,$
@@
$f_{T_{1}}(t;\Gamma_{1},a,b):$ $=\bar{F}(\rho(0),\rho(t))$ (11)
$=(1-a-b)+a\,e^{-\Gamma_{1}t/2}+b\,e^{-\Gamma_{1}t}\,,$
@@
$S_{ZZ}=\prod_{(i,j)\in G}\prod_{\Delta t\,\in\,\text{idle intervals{\hbox{(i,j)}}}}f_{ZZ}(\Delta t;(\omega_{ZZ})_{ij},c)\,.$ (15)
@@ Appendix B Loss functions
One important consideration of the pairwise loss, compared to the pointwise MSE loss, is that it facilitates a quadratic expansion of the dataset: an initial dataset of $N$ distinct circuits yields $N(N-1)/2$ distinct pairs of circuits. This expansion can be useful because it allows for many more training examples to be drawn from a given dataset, but it also has drawbacks because the dataset size can become unwieldy as well as depart from being independent and identically distributed (iid) [37]. For these reasons, we did not implement a pairwise loss in our experiments.
@@
Figure 5: Dataset statistics. (Top): The distribution of layout sizes (the number of ways that a given initial circuit can “fit” onto the device after hardware transpilation), according to algorithm. (Bottom): The distribution of empirical Hellinger fidelities, also according to algorithm.
@@ Appendix C Circuit dataset
In Fig. 5 we show the distribution of layouts (per initial circuit) according to algorithm, as well as the distribution of mean Hellinger fidelities for each initial circuit. The number of layouts per circuit ranges from 2 to 46. The details of each algorithm, such as the circuit width $n$ or the random graph used in the QAOA circuit were chosen to produce circuits whose Hellinger fidelities were not extreme.

### v2

@@ 1 Introduction
Extrapolating from hardware-level proxies to system-level performance is complicated by the fact that most common scalable QCVV approaches are by design insensitive to the myriad sub-structures, or contexts, that can appear within quantum circuits and which can have important impacts on the performance. This is in part due to the historic use of these metrics in analyses of fault tolerance, in which errors are typically (and speciously) assumed to be statistically independent. For example, randomized benchmarking (RB) is relatively efficient to implement and easy to interpret, yet it masks coherent errors exhibiting correlations in time over many gates in a circuit [10, 11]. Similarly, crosstalk may change both the magnitude and the nature of the error of a specific gate depending on the existence or absence of simultaneous gates on neighboring qubits [11, 12] - contexts that are ignored or averaged over in standard gate-level QCVV protocols. In particular, errors that occur during idling periods are in general poorly captured by most standard methods, both at the single qubit level where $T_{2}$ effects are relevant, and at the multi-qubit level where quantum crosstalk can lead to an unwanted unitary evolution between neighboring qubits [13, 14, 15]. Furthermore, the overall effect of an error on the measured fidelity of a quantum circuit will be dependent on the structure of the circuit; random circuits will generally be less sensitive to over-rotation gate errors due to self-cancellations of the random excess angles [16], whereas circuits with a repetitive structure, such as Trotter circuits, will be very sensitive to these types of errors [8, 10, 17]. The aggregate effect of these various circumstances - all of which are routinely encountered in the execution of real algorithms on hardware - tends to cause QCVV protocols to only be loosely predictive of algorithmic performance [9, 14].
Here, we focus on the concrete problem of layout selection within a single processor: the mapping of virtual or abstract qubits in a compiled circuit to physical device qubits in a manner that maximizes circuit fidelity. Previous work has examined the topic of layout selection via various methods [18, 19, 20, 21, 22, 23] (see also [24, 25]), but has not presented a comprehensive ranking procedure based on measurements of real circuits executed on hardware. This omission ignores the essential contextual information that is known to cause divergence between proxy measure predictions and achievable performance.
@@
In this work, we develop a new form of predictive QCVV for layout selection for arbitrary target circuits based on a heuristic, machine learning solution. We report direct measurements on an IBM quantum computer revealing that different layouts of individual target circuits exhibit up to a factor of $\times1$ variation in the ratio of the maximum achieved to median Hellinger fidelity, highlighting the importance of this task. We develop a procedure to rank the expected relative performance of logically equivalent layouts for arbitrary target circuits within a quantum processor, drawing from a large body of work in machine learning where this general problem is known as “Learning to Rank” [26, 27, 28, 29]. In so doing, this method prioritizes the relative performance between logically equivalent permutations of a user-defined target circuit over accurate estimates of absolute performance, and omits the reliance on human-interpretable proxy measures. Our method trains a circuit-score model over a dataset of circuits exhibiting diverse structural characteristics, and parameterizes the target-circuit score in a physics-based phenomenological manner. The model then assigns to each layout of the target circuit a numerical score used to quantify relative expected performance (a higher score corresponds to a higher expected performance). We test the efficacy of the ranking procedure using direct fidelity measurements of the different layouts on hardware, and compare our learning-to-rank method against alternative layout-selection approaches. We also compare five different candidate loss functions and quantitatively demonstrate that the listwise Rank MSE loss function performs best across multiple metrics. For the best model, measurements demonstrate a reduction in selection error – the relative loss in Hellinger fidelity due to sub-optimal layout selection – by 3.2×3.2$\times$3.2 × (1.8×1.8$\times$1.8 ×) when compared to a random and a baseline error-informed layout selection strategy, respectively, validating the utility of this approach.
The rest of the article is organized as follows. In Sec. 2 we briefly review the problem of layout selection. In Sec. 3 we introduce our approach, and then perform experiments on real quantum hardware in Sec. 4, before concluding with a discussion in Sec. 5. Further technical details are included in Appendix A and B.
2 The layout-selection problem
@@
3 Circuit ranking for layout selection
@@
We now present an overview of a practical and efficient machine-learning approach for ranking-based layout selection. We use a score-based approach for ranking circuit performance. The circuit score $S$ is a function that maps scheduled circuits to a real-valued numerical score: higher scores correspond to higher estimated fidelities. In a good ranking procedure, ordering layouts by circuit score should lead to the same ordinal ranking among circuits as a chosen performance metric derived from measurements on hardware. In this work, we take this measured quantity to be the Hellinger fidelity $H_{F}$ between the ideal bitstring distribution and the empirically observed bitstring distribution.
@@
$H_{F}(C_{1})\lesssim H_{F}(C_{2})\quad\text{iff}\quad S(C_{1})\lesssim S(C_{2})\,.$ (1)
Our objective is now to learn score functions that approximately obey Eq. 1.
@@
The loss function provides a continuous and differentiable way to penalize deviations from Eq. 1. We consider multiple loss functions, summarized in Table 1, reflecting the fact that there does not appear to be a clear preferred choice in the Learning-to-Rank literature. These can be grouped into three categories: pointwise, pairwise, and listwise. Pointwise approaches do not involve any comparisons, pairwise approaches involve comparisons between pairs of layouts, and listwise approaches involve comparisons across a set of all possible layouts for a given initial circuit. One important consideration of the pairwise loss, compared to the pointwise MSE loss, is that it facilitates a quadratic expansion of the dataset: an initial dataset of $N$ distinct circuits yields $N(N-1)/2$ distinct pairs of circuits. This expansion can be useful because it allows for many more training examples to be drawn from a given dataset, but it also has drawbacks because the dataset size can become unwieldy as well as depart from being independent and identically distributed (iid) [29]. For these reasons, we did not implement a pairwise loss in our experiments. Details for the various loss functions considered are presented in Appendix B.
@@
Score MSE pointwise Minimize difference between score and Hellinger fidelity
Pearson listwise Maximize correlation between score and Hellinger fidelity
Soft Spearman listwise Maximize correlation between score and Hellinger fidelity
Rank MSE listwise Minimize difference of true and predicted rank
NLL listwise Maximize model probability of observed top-$K$ layouts
@@
A summary of the machine learning pipeline for layout selection is presented in Fig. 3. To train the ML model, a dataset of circuits is first generated. This is illustrated in Fig. 3(a), where each layer represents a batch of layouts transpiled to the hardware for a different circuit. Each layout is then executed on the quantum processing unit (QPU) and the Hellinger fidelity is computed for all layouts by comparing the ideal bitstring distribution against the empirical distribution. Together, these batches of executed layouts and their fidelities form a dataset of labeled data. In the training phase, illustrated in Fig. 3(b), the batches of layouts in the dataset are ranked and the model iteratively updated. For each batch, the current state of the model produces a score ranking that is compared against the true ranking derived from the Hellinger fidelity. Mismatches in the two rankings are penalized by the differentiable loss $\mathcal{L}(\theta)$. The model parameters $\theta$ are updated through multiple iterations of a gradient-based algorithm. In the inference phase, illustrated in Fig. 3(c), the goal is to use the trained model to predict a single layout for execution on the QPU. The input is a single transpiled circuit and the most up-to-date backend information (such as gate errors and $T_{1}$ times). The set of all valid layouts is generated using the subgraph isomorphism capability of the Mapomatic package, and a score $S$ is assigned to each layout by the already trained model. The layouts are ranked according to the circuit score, and the circuit with the largest score is chosen for execution on the device.
4 Experimental results
@@
We benchmark our approach against both random selection and selection via the Mapomatic package [35]. Random selection is the default approach when utilizing a standard compiler. Many informed users employ Mapomatic, a layout selection method that estimates circuit fidelity by aggregating the backend-estimated fidelities of each individual operation (i.e., gate errors, measurement errors, etc). 222 Note that Mapomatic is the name of both the layout selection algorithm as well as the Python package that first identifies all valid layouts and then ranks them according to a simple heuristic score - here, we will use Mapomatic to specifically refer to the score-based layout selection algorithm. Mapomatic also uses a scoring function to rank layouts; the key points of difference are in the choice of score function and the fact that in the models developed here, the parameters are learned by training the model over a dataset containing a diverse range of circuit contexts.
@@
Table 2: Model performance. The performance of the circuit score model is measured by multiple evaluative metrics. Median rank $r$ corresponds to the median normed rank, Median rank $R$ corresponds to the median un-normalized rank, Selection Err corresponds to the selection error, and Top-1 Acc corresponds to the Top-1 accuracy. Each row corresponds to a circuit score model trained on a different loss function, averaged over 10 random seeds. Uncertainties correspond to 1 standard error. For comparison, the results for random layout selection and Mapomatic are also shown. The win rate of each method against Mapomatic (MM) is shown in the final column.
Model Loss Median rank $r$ Median rank $R$ Selection Err (%) Top-1 Acc (%) Win Rate (%)
@@
Rank MSEd=1 0.13275(39) 3.70(15) 10.603(97) 22.41(41) 67.19(18)
Rank MSEd=2 0.13275(39) 3.70(15) 10.603(97) 22.41(41) 67.19(18)
Rank MSEd=3 0.13304(29) 3.80(13) 10.638(78) 22.52(46) 67.03(18)
@@ 5 Discussion
Data availability
The data supporting the findings of this study are publicly available in Zenodo at https://doi.org/10.5281/zenodo.14037434 [40].
@@ Appendix A Phenomenological circuit score
where msmti represents the measurement operation on qubit $i$, and $f(\text{msmt}_{i})$ corresponds to one minus the mean miss-classification rate. Note that measurement error mitigation has been applied to the executed circuits, effectively reducing this error rate.
@@
$F(\rho(0),\rho(t))=|\alpha|^{2}$ $+\left(|\alpha|^{4}+|\beta|^{4}-|\alpha|^{2}\right)e^{-\Gamma_{1}t}+2|\alpha|^{2}|\beta|^{2}e^{-\Gamma_{2}t}\,,$ (10)
@@
$f_{T_{1}}(t;\Gamma_{1},a,b):$ $=\bar{F}(\rho(0),\rho(t))=(1-a-b)+a\,e^{-\Gamma_{1}t/2}+b\,e^{-\Gamma_{1}t}\,,$ (11)
@@
$S_{ZZ}=\prod_{(i,j)\in G}\prod_{\Delta t\,\in\,\text{idle intervals$(i,j)$}}f_{ZZ}(\Delta t;(\omega_{ZZ})_{ij},c)\,.$ (15)
@@ Appendix B Loss functions
As discussed in the main text, the pairwise loss necessitates a quadratic expansion of the dataset: an initial dataset of $N$ distinct circuits yields $N(N-1)/2$ distinct pairs of circuits. This expansion can be useful because it allows for many more training examples to be drawn from a given dataset, but it also has drawbacks because the dataset size can become unwieldy as well as depart from being independent and identically distributed (iid) [29]. For these reasons, we did not implement a pairwise loss in our experiments.
@@
Figure 5: Dataset statistics. (Top): The distribution of layout sizes (the number of ways that a given initial circuit can “fit” onto the device after hardware transpilation), according to algorithm. The number of layouts per circuit ranges from 2 to 46. (Bottom): The distribution of empirical Hellinger fidelities, also according to algorithm. Note that the details of each algorithm, such as the circuit width $n$ or the random graph used in the QAOA circuit were chosen to produce circuits whose Hellinger fidelities were not extreme (i.e., near 0 or near 1).
@@ Appendix C Circuit dataset
Lastly, in Fig. 5 we show the distribution of layouts (per initial circuit) according to algorithm, as well as the distribution of mean Hellinger fidelities for each initial circuit.